LIF STUDIES OF SIMPLE HALOCARBENES

\underline{H. FAN}, I. IONESCU, AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

We will report on new laser induced fluorescence (LIF) measurements of the A^{1}A′′-X^{1}A′ system in the monohalocarbenes HCCl and HCF, and the A^{1}B_{1}-X^{1}A_{1} system in the dihalocarbene CCl_{2}. The carbenes were generated under jet-cooled conditions using a pulsed discharge source. We have applied a variety of techniques to examine the excited state properties, and results from fluorescence excitation spectroscopy, single vibronic level lifetime measurements, and quantum beat spectroscopy will be reported.